I would like to know how to graph the greatest integer function [2x-3]=y.
please help

Answer by stanbon(48558)

I would like to know how to graph the greatest integer function [2x-3]=y.
-----------------------------------
Solve 2x-3 = 0
2x = 3
x= 3/2
=============
Mark the point (3/2,0) on a coordinate system.
--------------
Solve 2x-3 = 1
2x = 4
x = 2
---------------
Mark the point (2,1) on a coordinate system.
---------------
Draw a line segment on the x-axis from x=3/2 to x=2.
The left end has a solid point; the right end has an open circle.
--------------------
Draw a horizontal line segment at level "1" from x=2 to x=(5/2)
The left end has a solid point; the right end has an open circle.
---------------------
Continue this same pattern to the right and to the left.
Each line segment is "1/2" long.
The y-values increase by "1" for each segment.
The closed circle is always on the left end of the line segment.
The open circle is always on the right end of the line segment.
================
Cheers,
Stan H.